# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 21 Apr 2026 v4

@@ 7 Explicit Lower Bounds via the Number of Distinct Prime Factors
$c(p)\;\geq\;P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\qquad P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779,$
where $p_{1}<p_{2}<\cdots<p_{\omega}$ are the first $\omega$ primes.
@@
Since $P(t):=\prod_{j=1}^{\infty}(1-t^{-j})$ is increasing in $t>1$, we have $P(p)\geq P(3)=P_{3}$. Therefore, by (1),
@@
Corollary 7 (Matching lower bound).
@@
The lower bound follows by the same primorial/Mertens argument used in the proof of Corollary 4: for every odd prime $p\leq x$, $\omega(p-1)\leq m(x)$, and Lemma 1 together with Mertens’ third theorem gives $\min_{p\leq x,\,p\textup{ prime}}c(p)\gg 1/\!\log\log x$. Theorem 9 sharpens the constant from $\tfrac{1}{2}$ to $P_{3}$ but does not affect the asymptotic order. The matching upper bound is Proposition 1. ∎
Corollary 8 (Pointwise universal lower bound).
Let
@@
Finally, by Mertens’ third theorem and Lemma 2,
@@
Remark 4 (On effectivity).
The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully effective by combining explicit Mertens estimates (Rosser–Schoenfeld [34, Theorems 4 and 7], Dusart [12, Theorem 5.9]) with the Linnik-type bound $L\leq 5.18$ of Xylouris [41, Theorem 1.1].
Remark 5 (Sharper bounds for specific primes).
Direct computation of the Euler product gives $P(5)\approx 0.7603$, $P(101)\approx 0.9900$, and $P(p)>1-10^{-9}$ for $p>2^{30}$. For the NIST moduli $p=3329$ and $p=8380417$ [31, 32], $\prod_{j\geq 1}(1-p^{-j})>0.9996$ and >0.9999998>0.9999998, respectively.
Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$; by Theorem 9, $c(p)\geq B(\omega(p-1))$ for every prime $p\geq 3$. Table 1 lists representative values of $B(K)$ for small values of $K$.
@@
Table 1: Representative values of the explicit lower bound $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ appearing in Theorem 9. Thus every prime $p\geq 3$ with $\omega(p-1)\leq K$ satisfies $c(p)\geq B(K)$.
Remark 6 (On the case $\omega(p-1)=1$).
If $\omega(p-1)=1$ and $p$ is an odd prime, then p1p-1 is an even prime power, hence $p-1=2^{m}$ for some $m\geq 1$. Equivalently, $p$ is a Fermat prime. In this case
$\frac{\varphi(p-1)}{p-1}=\frac{1}{2},\qquad c(p)=\frac{1}{2}\,P(p),\qquad P(p):=\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
Hence $c(p)\to 1/2$ along any infinite subsequence of Fermat primes, so $B(1)$ is not tight for Fermat primes.
Corollary 9 (Explicit lower bounds for large primes).
@@
The claim under $\omega(p-1)\leq 10$ follows from the bound $c(p)\geq B(\omega(p-1))$ (Theorem 9), since
@@
By Dusart’s explicit product estimate [12, Theorem 5.9], which we use with threshold $x\geq 2\,278\,382$ (noting that $2\,278\,382<2^{22}\ll 2^{30}$, so the threshold is satisfied for all primes $p>2^{30}$),
$\prod_{\begin{subarray}{c}r\leq x\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr)\;=\;\frac{e^{-\gamma}}{\log x}\,\Bigl(1+O^{*}\!\Bigl(\frac{0.2}{\log^{3}x}\Bigr)\Bigr).$
Since $p>2^{30}$, we have p1>2 278 382p-1>2\,278\,382 and $\log(p-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
@@
$\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{p^{j}}=1-\frac{1}{p-1}>0.9999.$
@@
$c(p)=\frac{\varphi(p-1)}{p-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(p-1)}\Bigr)\cdot 0.9999\cdot\frac{e^{-\gamma}}{\log(p-1)}\;>\;\frac{0.56}{\log p},$
using $e^{-\gamma}>0.5614$, $\log(p-1)<\log p$, and $0.2/\log^{3}(p-1)<2.5\times 10^{-5}$ for $p>2^{30}$. Equivalently, for $p>2^{30}$,
@@
Remark 7.
By the shifted-prime Erdős–Kac theorem (Barban [3]; see also Halberstam [20] and Elliott [13, Ch. 12]), extending the classical result of Erdős and Kac [15], $(\omega(p-1)-\log\log x)/\sqrt{\log\log x}$ converges in distribution to the standard normal over primes $p\leq x$. For $p\approx 2^{128}$, $\log\log p\approx 4.5$, so $\omega(p-1)\leq 10$ holds for the vast majority of such primes.
Standard $p$ generic $B(\omega(p-1))$ bound sharper factorization bound
ML-KEM $3329$ $B(2)=P_{3}/3\approx 0.1867$ $\approx 0.4614$
ML-DSA $8380417$ $B(4)=8P_{3}/35\approx 0.1280$ $\approx 0.2933$
Table 2: The prime moduli appearing in FIPS 203 and FIPS 204 [31, 32]. The third column gives the generic bound $B(\omega(p-1))$ from Theorem 9, depending only on the number of distinct prime factors of p1p-1, while the fourth gives the sharper lower bound obtained from the exact factorization of p1p-1 together with $P(p)\geq 1-\frac{1}{p-1}$.
For the NIST-standardized moduli [31, 32], the exact factorizations $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$ together with $P(p)\geq 1-1/(p-1)$ give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$. The certified evaluation of Proposition 8 yields $c(3329)=0.461399\ldots$ and $c(8380417)=0.293255\ldots$, so the sharper factorization bounds agree with the exact values to $10$ and $21$ decimal places, respectively; the gap is $O(p^{-3})$, since replacing $P(p)$ by 11/(p1)1-1/(p-1) incurs an error of order $p^{-1}\cdot p^{-2}$ from the leading cross-term in the Euler product expansion.
Context $p$ $\omega(p-1)$ $1/c(p)$
ML-KEM (FIPS 203) $3329$ 22 $2.17$
ML-DSA (FIPS 204) $8380417$ 44 $3.41$
NewHope / FALCON [2, 19] $12289$ 22 $3.00$
BFV/BGV FHE (small) [5, 18, 6] $786433$ 22 $3.00$
BFV/BGV FHE (medium) [5, 18, 6] $2013265921$ 33 $3.75$
NTTRU / BLISS [27, 11] $7681$ 33 $3.75$
64-bit NTT prime [37] $2^{64}-2^{32}+1$ 66 $4.00$
Fermat prime ($2^{16}+1$) $65537$ 11 $2.00$
Table 3: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(p-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(p)$. The displayed overheads are rounded to two decimal places; the exact values differ by less than $3\times 10^{-3}$.
Remark 8 (Concrete example).
For the 64-bit NTT prime $p=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime”; not to be confused with the Ed448-Goldilocks prime [21]), used in lattice cryptography [37], one has $\omega(p-1)=6$, $\varphi(p-1)/(p-1)\approx 0.2500$, $P(p)>1-10^{-19}$, and $1/c(p)\approx 4$.
Proposition 8 (Certified evaluation of $c(p)$).
Given a prime $p\geq 3$ and the complete factorization $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$, the following procedure computes $c(p)$ to $D$ correct decimal digits using $O(k+D\log 10/\!\log p)$ rational-arithmetic operations in the unit-cost rational-arithmetic model:
@@
8 Concluding Remarks
This article studies the PRIM-LWE density constant $c(p)$ unconditionally from extremal, distributional, and computational viewpoints. We proved that $\inf_{p\text{ prime}}c(p)=0$ and that $\min_{p\leq x}c(p)\asymp 1/\log\log x$, so the worst-case rejection-sampling overhead grows only on a doubly logarithmic scale. We also showed that $c(p)$ admits a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$ and Hausdorff dimension 0; in particular, the limiting measure is supported on a metrically negligible subset of $0,\tfrac{1}{2}]$ despite having full topological support. Finally, the explicit $\omega(p-1)$-based bounds and the certified evaluation algorithm of Proposition 8 make these asymptotic results effective for concrete cryptographic moduli.
Several questions remain open. The squared modulus of the characteristic function admits an explicit Euler product (Proposition 5), but whether

### 06 May 2026 v5

@@ 7. Explicit Lower Bounds via the Number of Distinct Prime Factors
$c(p)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\>P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779$
(that is, $P_{3}=P(3)$), where $p_{1}<p_{2}<\cdots<p_{\omega}$ are the first $\omega$ primes.
@@
Since $P(t):=\prod_{j=1}^{\infty}(1-t^{-j})$ is increasing in $t>1$ (each factor $1-t^{-j}$ is positive and increasing for $t>1$, and the product converges), we have $P(p)\geq P(3)=P_{3}$. Therefore, by (2.1),
@@
Corollary 7.2 (Matching lower bound).
@@
The prime $p=2$ satisfies $c(2)=\prod_{j\geq 1}(1-2^{-j})\approx 0.2888$, a fixed positive constant that exceeds $c/\!\log\log x$ for all large $x$. For every odd prime $p\leq x$, the primorial bound $N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1<x$ gives $\omega(p-1)\leq m(x)$. By Theorem 7.1 and the fact that each factor $(1-1/p_{i})\in(0,1)$,
$c(p)\;\geq\;P_{3}\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\geq\;P_{3}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Mertens’ third theorem and Lemma 4.1 ($p_{m(x)}\sim\log x$) now yield
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\geq\;P_{3}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\asymp\;\frac{1}{\log\log x}.$
The matching upper bound is Proposition 4.2. ∎
Corollary 7.3 (Pointwise universal lower bound).
Recall the primorial-counting function from Section 4:
@@
Finally, by Mertens’ third theorem and Lemma 4.1,
@@
Theorem 7.4 (Sharp constant for the extremal order).
$\limsup_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}\frac{1}{c(p)\,\log\log p}\;=\;e^{\gamma}.$
In particular, $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ for every prime $p$, and this leading coefficient is best possible.
Proof.
Upper bound. Let $p\geq 3$ be prime, with $\omega:=\omega(p-1)$ distinct prime factors $r_{1}<\cdots<r_{\omega}$ of p1p-1. Since $t\mapsto t/(t-1)$ is strictly decreasing for $t>1$ and $r_{i}\geq p_{i}$ for each $i$,
$\frac{p-1}{\varphi(p-1)}=\prod_{i=1}^{\omega}\frac{r_{i}}{r_{i}-1}\;\leq\;\prod_{i=1}^{\omega}\frac{p_{i}}{p_{i}-1}.$
The primorial bound $N_{\omega}=\prod_{i=1}^{\omega}p_{i}\leq\prod_{r\mid(p-1)}r\leq p-1$ gives $\omega\leq m(p-1)$, so extending the product (each factor exceeds $1$):
$\frac{p-1}{\varphi(p-1)}\;\leq\;\prod_{i=1}^{m(p-1)}\frac{p_{i}}{p_{i}-1}.$
By Mertens’ third theorem (Theorem 3.1) and Lemma 4.1,
$\prod_{i=1}^{m(p-1)}\frac{p_{i}}{p_{i}-1}\;\sim\;e^{\gamma}\log p_{m(p-1)}\;\sim\;e^{\gamma}\log\log p.$
Since $1/P(p)=1+O(1/p)\to 1$, we obtain $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$.
Lower bound. For each $k\geq 1$, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$, which exists by Dirichlet’s theorem. By Linnik’s theorem, $p^{*}(k)\leq C_{0}N_{k}^{\,L}$ for absolute constants $C_{0},L>0$.
Since $N_{k}\mid(p^{*}(k)-1)$, every prime $p_{i}$ with $i\leq k$ divides $p^{*}(k)-1$, and additional prime factors of $p^{*}(k)-1$ beyond $p_{k}$ only increase the product. Hence
$\frac{p^{*}(k)-1}{\varphi(p^{*}(k)-1)}\;\geq\;\prod_{i=1}^{k}\frac{p_{i}}{p_{i}-1}\;\sim\;e^{\gamma}\log p_{k}.$
For the asymptotics of $\log\log p^{*}(k)$: by Linnik’s bound, $\log p^{*}(k)\leq\log C_{0}+L\log N_{k}\sim Lp_{k}$ (using $\log N_{k}=\vartheta(p_{k})\sim p_{k}$ from Lemma 4.1), so $\log\log p^{*}(k)\leq\log p_{k}+\log L+o(1)$. Since $p^{*}(k)>N_{k}$ and $\log N_{k}\sim p_{k}$, we also have $\log\log p^{*}(k)\geq\log p_{k}+o(1)$. Therefore $\log\log p^{*}(k)\sim\log p_{k}$. Combined with $P(p^{*}(k))\to 1$:
$\frac{1}{c(p^{*}(k))\,\log\log p^{*}(k)}\;\geq\;\frac{(1+o(1))\,e^{\gamma}\log p_{k}}{(1+o(1))\,\log p_{k}}\;\to\;e^{\gamma}.$
Together, the two bounds yield $\limsup=e^{\gamma}$. ∎
Remark 7.5 (On effectivity).
The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully effective by combining explicit Mertens estimates (Rosser–Schoenfeld [29, Theorems 4 and 7], Dusart [9, Theorem 5.9]) with the Linnik-type bound $L\leq 5.18$ of Xylouris [37, Theorem 1.1], whose implied constant is effectively computable (though no numerical value has been published).
Remark 7.6 (Convergence rate of the Euler product).
Direct computation gives $P(5)\approx 0.7603$, $P(101)\approx 0.9900$, and $P(p)>1-10^{-9}$ for $p>2^{30}$; the correction factor $P(p)$ is thus negligible for all but the smallest primes.
Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$; by Theorem 7.1, $c(p)\geq B(\omega(p-1))$ for every prime $p\geq 3$. Table 1 lists representative values of $B(K)$ for small values of $K$.
@@
Table 1. Representative values of the explicit lower bound $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ appearing in Theorem 7.1. Thus every prime $p\geq 3$ with $\omega(p-1)\leq K$ satisfies $c(p)\geq B(K)$.
Remark 7.7 (On the case $\omega(p-1)=1$).
If $\omega(p-1)=1$ then $p$ is a Fermat prime, $\varphi(p-1)/(p-1)=1/2$, and $c(p)=P(p)/2$; since $P(p)>P(3)$ for every prime $p\geq 5$, each known Fermat prime $p\geq 5$ satisfies $c(p)>B(1)$, so the bound is not tight. Along any unbounded sequence of Fermat primes, should one exist, $P(p)\to 1$ and thus $c(p)\to 1/2$.
Corollary 7.8 (Fully explicit coarse lower bound for large primes).
@@
The claim under $\omega(p-1)\leq 10$ follows from the bound $c(p)\geq B(\omega(p-1))$ (Theorem 7.1), since
@@
By Dusart’s explicit product estimate [9, Theorem 5.9], which we use with threshold $x\geq 2\,278\,382$ (noting that $2\,278\,382<2^{22}\ll 2^{30}$, so the threshold is satisfied for all primes $p>2^{30}$),
$\prod_{\begin{subarray}{c}r\leq x\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr)\;=\;\frac{e^{-\gamma}}{\log x}\,\Bigl(1+O^{*}\!\Bigl(\frac{0.2}{\log^{3}x}\Bigr)\Bigr),$
where $O^{*}(\delta)$ denotes a quantity of absolute value at most $\delta$. Since $p>2^{30}$, we have p1>2 278 382p-1>2\,278\,382 and $\log(p-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
@@
$\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{p^{j}}=1-\frac{1}{p-1}>1-10^{-9}.$
@@
$c(p)$ $=\frac{\varphi(p-1)}{p-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(p-1)}\Bigr)(1-10^{-9})\,\frac{e^{-\gamma}}{\log(p-1)}$
$>\;\frac{0.56}{\log p},$
using $e^{-\gamma}>0.5614$, $\log(p-1)<\log p$, $0.2/\log^{3}(p-1)<2.5\times 10^{-5}$, and $1/(p-1)<10^{-9}$ for $p>2^{30}$. Equivalently, for $p>2^{30}$,
@@
The $1/c(p)<1.79\log p$ bound is explicit and valid uniformly for every prime $p>2^{30}$ without any factorisation data. The sharp asymptotic scale for the extremal order is the much slower $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ of Theorem 7.4.
Remark 7.9 (Evaluation of $c(p)$).
Given the factorisation of p1p-1, the value $c(p)$ can be computed to $D$ correct decimal places (absolute error < $10^{-D}$) using $O(k+D\log 10/\!\log p)$ rational-arithmetic operations, where $k=\omega(p-1)$ is the number of distinct prime factors of p1p-1, by computing the totient ratio $\varphi(p-1)/(p-1)$ exactly and truncating the Euler product $P(p)$ at depth $J=\lceil D\log 10/\!\log p\rceil$; writing $P_{J}(p):=\prod_{j=1}^{J}(1-p^{-j})$, the tail satisfies $0\leq P_{J}(p)-P(p)\leq p^{-J}/(p-1)$.
8. Cryptographic Consequences
The density $c(p)$ arises directly as a reduction constant in the PRIM-LWE variant of the learning with errors (LWE) problem [47], introduced by Sehrawat, Yeo, and Desmedt [31] to construct verifiable secret sharing schemes with “free” verification. The LWE-to-PRIM-LWE reduction proceeds by rejection-sampling: a uniformly random $n\times n$ matrix over $\mathbb{F}_{p}$ is accepted if and only if its determinant is a primitive root modulo $p$, and the acceptance probability in dimension $n$ is $c_{n}(p)$ (see Section 2). Since $c_{n}(p)$ is decreasing in $n$ with limit $c(p)$, the dimension-uniform expected rejection-sampling overhead is $1/c(p)$. This overhead represents a tightness loss in the security reduction: by the analysis of [31], the hardness guarantee for PRIM-LWE degrades by a factor of $c(p)$ relative to standard LWE, making tight bounds on $c(p)$ essential for concrete security estimates. Since $\inf_{p\textup{ prime}}c(p)=0$ (Theorem 3.3), the overhead is unbounded in general; the explicit bounds of Section 7 quantify it for any given prime.
The results of the preceding sections translate immediately into cryptographic bounds. We collect the principal consequences.
Corollary 8.1 (Compact overhead bound).
Let $P_{3}$ be as in Theorem 7.1. If $p\geq 3$ is a prime modulus with $\omega(p-1)\leq K$, then the dimension-uniform expected overhead in the LWE-to-PRIM-LWE reduction satisfies
$\frac{1}{c(p)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
Proof.
Immediate from Theorem 7.1 upon taking reciprocals. ∎
For concrete deployment, it is useful to evaluate $c(p)$ exactly. The following proposition upgrades the evaluation procedure described in Remark 7.9 to a certified algorithm with explicit error control.
Proposition 8.2 (Certified evaluation of $c(p)$).
Given a prime $p\geq 3$ and the complete factorisation $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$, the following procedure computes $c(p)$ to $D$ correct decimal digits using $O(k+D\log 10/\!\log p)$ rational-arithmetic operations in the unit-cost rational-arithmetic model:
@@ 8. Cryptographic Consequences
Proposition 8.2 and the exact factorisations of p1p-1 for the NIST-standardised moduli [45, 46] yield the bounds recorded in Table 2. The factorisations $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$, together with $P(p)\geq 1-1/(p-1)$, give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$. The certified evaluation yields $c(3329)=0.461399\ldots$ and $c(8380417)=0.293255\ldots$, so the sharper factorisation bounds agree with the exact values to $10$ and $21$ decimal places, respectively; the gap is $O(p^{-3})$, since replacing $P(p)$ by 11/(p1)1-1/(p-1) incurs an error of order $p^{-1}\cdot p^{-2}$ from the leading cross-term in the Euler product expansion.
Standard $p$ generic $B(\omega(p-1))$ bound sharper factorisation bound
ML-KEM (FIPS 203) $3329$ $B(2)=P_{3}/3\approx 0.1867$ $\approx 0.4614$
ML-DSA (FIPS 204) $8380417$ $B(4)=8P_{3}/35\approx 0.1280$ $\approx 0.2933$
Table 2. The prime moduli appearing in FIPS 203 and FIPS 204 [45, 46]. The third column gives the generic bound $B(\omega(p-1))$ from Theorem 7.1, depending only on the number of distinct prime factors of p1p-1, while the fourth gives the sharper lower bound obtained from the exact factorisation of p1p-1 together with $P(p)\geq 1-1/(p-1)$.
Table 3 collects the overhead $1/c(p)$ for representative NTT-friendly primes from lattice-based cryptography and fully homomorphic encryption implementations.
Context $p$ $\omega(p-1)$ $1/c(p)$
ML-KEM (FIPS 203) $3329$ 22 $2.17$
ML-DSA (FIPS 204) $8380417$ 44 $3.41$
NewHope / FALCON [38, 42] $12289$ 22 $3.00$
BFV/BGV FHE (small) [40, 43, 39] $786433$ 22 $3.00$
BFV/BGV FHE (medium) [40, 43, 39] $2013265921$ 33 $3.75$
NTTRU / BLISS [44, 41] $7681$ 33 $3.75$
64-bit NTT prime [48] $2^{64}-2^{32}+1$ 66 $4.00$
Fermat prime ($2^{16}+1$) $65537$ 11 $2.00$
Table 3. Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(p-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(p)$. The displayed overheads are rounded to two decimal places; the exact values differ by less than $3\times 10^{-3}$.
Remark 8.3 (Typical overhead at cryptographic scales).
By the shifted-prime Erdős–Kac theorem (Section 7), $\omega(p-1)$ is typically of order $\log\log p$. For $p\approx 2^{128}$, $\log\log p\approx 4.5$, so $\omega(p-1)\leq 10$ holds for the vast majority of such primes, and Table 1 gives $c(p)>0.088$. In practice, Table 3 shows $1/c(p)\leq 4$ for all surveyed moduli. By Corollary 7.8, the overhead satisfies $1/c(p)<1.79\log p$ uniformly for every prime $p>2^{30}$ without any factorisation data; the sharp asymptotic scale is $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ (Theorem 7.4), and the order $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ is tight (Corollary 7.2).
Remark 8.4 (64-bit NTT prime).
For the 64-bit NTT prime $p=2^{64}-2^{32}+1$ used in lattice cryptography [48], one has $\omega(p-1)=6$, $\varphi(p-1)/(p-1)\approx 0.2500$, $P(p)>1-10^{-19}$, and $1/c(p)\approx 4$.
9. Concluding Remarks and Open Questions
This article studies the primitive-root determinant density $c(p)$ from extremal, distributional, and metric viewpoints. We proved that $\inf_{p\text{ prime}}c(p)=0$ and that $\min_{p\leq x}c(p)\asymp 1/\log\log x$, so the extremal order of the reciprocal grows only on a doubly logarithmic scale, with sharp leading coefficient $e^{\gamma}$. We also showed that $c(p)$ admits a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$ and Hausdorff dimension 0; in particular, the limiting measure is supported on a Lebesgue-null subset of $0,\tfrac{1}{2}]$ despite having full topological support. The explicit $\omega(p-1)$-based bounds make these asymptotic results effective. Section 8 translates the explicit bounds into concrete overhead estimates for the PRIM-LWE reduction in lattice-based cryptography.
We conclude with two open questions.
1. Fourier decay and the Rajchman property. The squared modulus of the characteristic function admits an explicit Euler product (Proposition 6.6), but whether